1. (6 points) Draw each angle in standard position, then find and draw a coterminal angel. 371 177
Answer #1

-SAN 9x/+ -64)२ -N/A, A/4 roterminal of o= 37 Be+ 360 -185+360 २२5. 1 S/ 1-3114 -714 -२०/4 be = -1+x/6 -44/6 -87/ -1076 - -11